OpenStudy (anonymous):

A circle has a diameter of 15 feet. What is the circumference of the circle? (Express your answer using π.)

OpenStudy (anonymous):

(Express your answer using pi

OpenStudy (anonymous):

@conain007 @AnswerMyQuestions

OpenStudy (anonymous):

Same formula.

OpenStudy (anonymous):

2 x pi x radius.

OpenStudy (anonymous):

But it gives us the diameter. And the radius is half the diameter, so what is your answer?

OpenStudy (anonymous):

15 / 2 =7.5

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So what is the circumference?

OpenStudy (anonymous):

47.1 !

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

but thats not any of the choices :p

OpenStudy (anonymous):

Hmm, what are your choices?

OpenStudy (anonymous):

The formula for the circumference of a circle is: \(C = 2 \pi r\) The diameter is 15, so the radius is 7.5. \(C = 2 \pi (7.5)\) What's 7.5 * 2?

OpenStudy (anonymous):

Can you multiply that?

OpenStudy (anonymous):

pi/15 feet 15pi feet 15/pi feet 30pi feet

OpenStudy (anonymous):

It gives us the same answer we ended up with.

OpenStudy (anonymous):

Multiply 7.5 * 2 and that is your answer in pi

OpenStudy (anonymous):

It would be 15 again.

OpenStudy (anonymous):

Oh..The choices are like that... :P

OpenStudy (anonymous):

Yes, so the answer is \(15 \pi\).

OpenStudy (anonymous):

@iGreen. and @AnswerMyQuestions Thanks :)

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

No problem. :)